OpenStudy (asad786):

Translate the sentence into an equation. Four times the difference of a number and 5 is 24. A. 4 • (c – 5) = 24 B. 4c – 5 = 24 C. c – 4 • 5 = 24 D. 5 • (c – 4) = 24

OpenStudy (asad786):

@studygurl

OpenStudy (asad786):

@studygurl14

OpenStudy (luffingsails):

"The difference of a number c and 5" = c - 5 "Four times the above" = 4 * (c - 5) = 24

OpenStudy (asad786):

a